Q20. Genovia has experienced exceptional growth in recent years. Its GDP per capita has increased from around $30,000 to $50,000 in last 5 years. Over the period quantity demanded of personal cars has increased from 450,000 units per year to 600,000 units. Quantity demanded of public transport, however, has declined from 10,000 buses to 7,000 buses. Calculate income elasticity of demand and tell which product is a normal good and which one is inferior.
1
Expert's answer
2019-09-04T09:07:47-0400

Income elasticity of demand for cars is:

"Ed = (600,000 - 450,000)\/(50,000 - 30,000)\u00d7(50,000 + 30,000)\/(600,000 + 450,000) = 150,000\/20,000\u00d780,000\/1,050,000 = 4\/7"so car is a normal good.

Income elasticity of demand for buses is:

"Ed = (7,000 - 10,000)\/(50,000 - 30,000)\u00d7(50,000 + 30,000)\/(7,000 + 10,000) = -3,000\/20,000\u00d780,000\/17,000 = - 12\/17,"

so public transport is an inferior good.